OpenStudy (anonymous):

I NEED TO SLEEP PLEASE ASSIST FOR STUDYING :) How do I simplify radical 5/6

OpenStudy (anonymous):

\[\sqrt{\frac{ 5 }{ 6 }}\]

OpenStudy (anonymous):

\[\sqrt{\frac{5}{6}}=\frac{\sqrt5}{\sqrt6}\] \[=\frac{\sqrt5}{\sqrt6}\times \frac{\sqrt6}{\sqrt6}\]

OpenStudy (anonymous):

when you multiply, the radical in the denominator will be gone, right? because \(\sqrt6\times \sqrt6=6\)

OpenStudy (anonymous):

OH yesss

OpenStudy (anonymous):

you will be left only with a radical in the numerator, namely \(\sqrt5\times \sqrt6=\sqrt{30}\)

OpenStudy (anonymous):

that wasn't too hard was it?

OpenStudy (anonymous):

ok well i get how to eliminate the radical from the denominator but that wat do I do with the other half

OpenStudy (anonymous):

you are done final answer is \[\frac{\sqrt{30}}{6}\]

OpenStudy (anonymous):

you can do no more over out halas get some sleep

OpenStudy (anonymous):

I understand now thank you so muchh

OpenStudy (anonymous):

I must continue so I will do good on the test tomorrow morning

OpenStudy (anonymous):

I need to understand the material

OpenStudy (anonymous):

ok you got another one?

OpenStudy (anonymous):

Ok well the next one i understand because of how you explained this problem thankfully but i do have another let me write it out

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[\left(\begin{matrix}3\sqrt{6} \\ 2\end{matrix}\right)2\] There should be a lne separating two as a denominator but you get it haha

OpenStudy (anonymous):

\[\left(\frac{3\sqrt{6}}{2}\right)^2\]?

OpenStudy (anonymous):

that is the same as \[\left(\frac{3\sqrt6}{2}\right)\times \left(\frac{3\sqrt6}{2}\right)\] multiply straight across

OpenStudy (anonymous):

the denominator is pretty obviously 4 right?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

would it be \[\frac{ 18 }{ 2 }\]

OpenStudy (anonymous):

how about the numerator? it is \[3\sqrt6\times 3\sqrt6\]

OpenStudy (anonymous):

The radical gets eliminated and the product equals whats inside

OpenStudy (anonymous):

oops \[3\times 3\times \sqrt6\times \sqrt6=9\times 6=54\]

OpenStudy (anonymous):

and \[\frac{54}{4}=\frac{27}{2}\] i think

OpenStudy (anonymous):

Yess you're right I think I messed you up the first time!

OpenStudy (anonymous):

I forgut to multiply the 3s

OpenStudy (anonymous):

i think maybe you forgot to square the 3

OpenStudy (anonymous):

Yea exacylu

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Thank you so much for all the help I'm finally complete with studying and ACTUALLY UNDERSTAND :))))

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

Goodnight!

OpenStudy (anonymous):

don't make any silly mistakes on the test good luck good night

OpenStudy (anonymous):

Thank you I will be sure of that